if total is $6.10 and there is a combination of nickles and dimes that add up to a total of 80 coins hom many nickles and how many dimes each are there
Answer by mananth(16949)

Nickels x
Dimes y

1 x + 1 y = 80 .............1
5 x + 10 y = 610 .............2

multiply (1)by -10
Multiply (2) by 1
-10 x -10 y = -800
5 x 10 y = 610
Add the two equations
-5 x = -190
/ -5
x = 38 Nickels
plug value of x in (1)
1 x + 1 y = 80
38 + 1 y = 80
1 y = 80 -38
1 y = 42
y = 42 Dimes